OpenStudy (sette):

Solve for k . i keep getting 11.1333333333 -20-6k=16-9(-k+19) k= ?

OpenStudy (anonymous):

k=9

OpenStudy (sette):

thanks

jimthompson5910 (jim_thompson5910):

-20-6k=16-9(-k+19) -20-6k=16+9k-171 -20-6k=9k-155 -6k=9k-155+20 -6k-9k=-155+20 -15k=-155+20 -15k=-135 k=-135/(-15) k=9